# Complex Valued Deep Operator Network (DeepONet) $[\mathcal{G}]$ for Three Dimensional Maxwell’s Equations: $\mathcal{G}\in\mathbb{C}^{m\times n}$

### v1

@@ 3.3 Imposing the PEC boundary condition
To impose the physics on the proposed DeepONet architecture, we adopt the approach of soft constraint. Since, the metallic scatterer examples studied in this work are perfect electric conductors (PEC), and the electric field on the surface of a PEC scatterer is orthogonal to the conductor. Therefore, the PEC boundary condition entails that $\bm{n}\times\bm{E}=0$, where $\bm{n}$ is the normal vector to the scatterer surface and $\boldsymbol{E}$ is the total field. Previous research such as [54, 55] has proposed ways of integrating such first-principle-informed conditions directly into the architecture of neural networks. Due to the high-dimensional training data in our problem, we instead incorporate the PEC boundary condition directly into the loss function.
Since a separate complex-valued DeepONet is trained for each field component, the boundary condition is enforced in two stages, as illustrated in Figure 3. First, DeepONets are trained independently and concurrently for the components $E_{x}$, $E_{y}$, $E_{z}$, $H_{x}$, $H_{y}$, and $H_{z}$, with each model minimizing the data-driven loss as defined in Equation 14. Once the data-driven loss for each component is minimized sufficiently, the trained models are saved. Next, to impose the boundary condition on the electric field, previously trained models for $E_{x}$, $E_{y}$, and $E_{z}$ are loaded and retrained jointly. During this phase of training, the PEC boundary condition is introduced into the loss function as an additional term. The loss function for this retraining step is defined as
$L=w_{1}||\bm{n}\times\bm{E}^{NN}||_{2}^{2}+w_{2}\left(\text{MSE}(E_{x}^{NN},\bar{E}_{x})+\text{MSE}(E_{y}^{NN},\bar{E}_{y})+\text{MSE}(E_{z}^{NN},\bar{E}_{z})\right),$ (15)
@@
Figure 3: Illustration of the two-stage process for training a combination of DeepONet models for different field components to impose the PEC boundary condition $\bm{n}\times\bm{E}=0$ on the electric field.
@@ 3.4 Reparametrization of the Complex-valued DeepONet
Specifically, let $\bm{u}=\{u_{k}\}_{k=1}^{K}$ be a set of complex-valued input functions from the function space $\mathcal{U}$, and let $v_{k}(\cdot)=\mathcal{G}[u_{k}](\cdot)$ be the corresponding complex-valued output functions in the space $\mathcal{V}$. The branch network, denoted as $\bm{b}(\cdot;\theta)$, is a CVNN with $L_{b}$ layers and an output dimension of $N$
@@ 3.5 Sampling of the input space
The performance of the DeepONet is sensitive to the values of the angular frequency selected for training. However, sampling a broad range of frequencies is inefficient and impractical due to the computational expense of generating large training datasets. Thus, an efficient data acquisition strategy is required to select frequencies that maximize information gain and enhance model performance with fewer samples. Traditional methods for sampling training frequencies within a given interval $[f_{min},f_{max}]$ such as uniform sampling, Latin hypercube sampling, and random sampling do not account for the sensitivity of the model’s predictions to input frequencies. Instead, we employ the method of acquisition function [57] to sample input space for the training of complex-valued DeepONet.
@@
4 Results
In this section we show the performance of the proposed DeepONet architecture for the Maxwell’s equation.
4.1 A comparison between Real valued DeepONet, Complex value DeepONet and Reparametrized complex valued DeepONet
@@ 4 Results
To demonstrate the superior performance of the complex-valued DeepONet compared to the vanilla real-valued DeepONet, we used similar architectures for both models, training them on the same 80 frequencies and testing on 21 frequencies sampled from a range of $[0.05,0.6]\,\text{GHz}$. The hyperparameters for both models used in our experiments are summarized in Table 1. For the real-valued DeepONet, the real and imaginary parts of a function are trained separately and then combined into a complex number for comparison against the ground-truth data. The training loss histories of the two models are plotted in Figure 4, with the dotted line representing the real-valued DeepONet and the solid line representing the complex-valued DeepONet. The loss stabilizes at around 10,000 epochs, with the complex-valued DeepONet achieving superior training loss across all components of the field. The final testing results are summarized in Figure 4, which shows that the reparametrized complex-valued DeepONet, trained using the two-step method, exhibits the best testing performance. The complex-valued models significantly outperform the vanilla real-valued model by decreasing the error with a factor ranging from 5-7 times. In terms of inference speed, the UWVF numerical solver takes 15 minutes to predict one frequency using 192 cores. In comparison, the complex-valued DeepONet predicts a single frequency in 1.16 seconds on a 24-core CPU or 0.10 seconds on an Nvidia Titan RTX GPU.
@@
Figure 5 shows the correlation between predicted and true total fields. Strong diagonal alignment demonstrates high prediction accuracy for both the electric and magnetic fields. Figure 6 shows the predictions of the total electric and magnetic fields by the complex-valued model at a testing frequency of 0.402 GHz. The pointwise absolute errors between the ground truths and predictions are plotted in green, with small errors primarily occurring in areas where the regularity of the function is poor. The model is shown to generalize well to the testing frequencies.
@@
Figure 4: Left sub figure shows training loss history for each electromagnetic field using complex-valued DeepONets (solid line) and real-valued DeepONets (dotted line), starting from 500 epochs. This clearly shows the improvement of the accuracy for complex valued DeepONet significantly. The right subfigure shows Testing performance of DeepONets (the complex-valued DeepONet, and the reparametrized complex-valued DeepONet) at testing frequencies. The plot infer a superior performance of complex valued DeepONet over real valued DeepONet.
@@
Figure 6: A comparison between ground truth (red solid line) and prediction (blue dotted line) of each component of the electric and magnetic fields at a frequency of 0.402 GHz. The pointwise absolute error between the true and predicted values is plotted by the green line. Figures show a very good agreement between actual and inferred electromagnetic field.
[…]
Component Training $L_{2}$ Testing $L_{2}$
@@
4.2 Acquisition functions for sampling the input space
@@
The acquisition function is used to efficiently select training frequencies for the complex-valued DeepONet. We consider the frequency range of $[0.05,0.6]\,\text{GHz}$. An initial set of 5 uniformly distributed frequencies in this range are selected for training an ensemble of 3 DeepONets. Then, an additional 15 frequencies are acquired to maximize information gain as determined by the acquisition function. The trained model is subsequently tested on 80 unseen samples.
@@
4.3 Deponent as surrogate for computing radar cross section (RCS) results
@@
The RCS is expressed as the ratio of the power density of the scattered field evaluated in the far-field to the incident power density times the area of a sphere of radius $r$, in the limit $r\to\infty$. The monostatic and bistatic radar cross sections (RCS), computed using field predictions from the complex-valued DeepONet, are presented in Figure 9. For these results, the DeepONet was trained on 60 samples and tested on 40 samples. In Figure 9, the monostatic RCS is calculated at an azimuth angle of $\phi=0$ for frequencies ranging from 0.05 to 0.6 GHz, while the bistatic RCS is shown for a frequency of $0.281$ GHz with $\phi$ varying from 00 to $360$ degrees. For both monostatic and bistatic RCS calculations, the analytical Mie-series solutions are used as the ground truth for comparison. The results of RCS ensures the accuracy proposed DeepONet model for computing the integrated quantity, which is very sensitive to the accuracy of predict electromagnetic field from DeepONets.
Figure 9: Monostatic and bistatic RCS results computed from DeepONet predictions (green) and compared to ground-truth Mie-series solutions (red). The monostatic RCS (left) is evaluated at 40 testing frequencies from the interval $[0.05,0.6]$ GHz. The bistatic RCS (right) is calculated at a single frequency of $0.281$ GHz and plotted against the azimuth angle $\phi$.

### v2

@@ 3.3 Imposing the PEC boundary condition
To impose the physics on the proposed DeepONet architecture, we adopt the approach of soft constraint. Since the metallic scatterer examples studied in this work are perfect electric conductors (PEC), the electric field on the surface of a PEC scatterer is orthogonal to the conductor. Therefore, the PEC boundary condition entails that $\bm{\hat{n}}\times\bm{E}=0$, where $\bm{\hat{n}}$ is the normal vector to the scatterer surface and $\boldsymbol{E}$ is the total field. Previous research such as [35, 36] has proposed ways of integrating such first-principle-informed conditions directly into the architecture of neural networks. Due to the high-dimensional training data in our problem, we instead incorporate the PEC boundary condition directly into the loss function.
Since a separate complex-valued DeepONet is trained for each field component, the boundary condition is enforced in two stages, as illustrated in Figure 2. First, DeepONets are trained independently and concurrently for the components $E_{x}$, $E_{y}$, $E_{z}$, $H_{x}$, $H_{y}$, and $H_{z}$, with each model minimizing the data-driven loss as defined in (13). Once the data-driven loss for each component is minimized sufficiently, the trained models are saved. Next, to impose the boundary condition on the electric field, previously trained models for $E_{x}$, $E_{y}$, and $E_{z}$ are loaded and retrained jointly. During this phase of training, the PEC boundary condition is introduced into the loss function as an additional term. The loss function for this retraining step is defined as
[…]
@@
Figure 2: Illustration of the two-stage process for training a combination of DeepONet models for different field components to impose the PEC boundary condition $\bm{\hat{n}}\times\bm{E}=0$ on the electric field.
@@ 3.4 Reparametrization of the Complex-valued DeepONet
Specifically, let $\bm{u}=\{u_{k}\}_{k=1}^{K}$ be a set of complex-valued input functions from the function space $\mathcal{U}$, and let $v_{k}(\cdot)=\mathcal{G}[u_{k}](\cdot)$ be the corresponding complex-valued output functions in the space $\mathcal{V}$. The branch network, denoted as $\bm{b}(\cdot;\theta)$, is a CVNN with $L_{b}$ layers and an output dimension of NN
@@ 3.5 Sampling of the input space
The performance of the DeepONet is sensitive to the values of the angular frequency selected for training. However, sampling a broad range of frequencies is inefficient and impractical due to the computational expense of generating large training datasets. Thus, an efficient data acquisition strategy is required to select frequencies that maximize information gain and enhance model performance with fewer samples. Traditional methods for sampling training frequencies within a given interval $f_{min},f_{max}]$ such as uniform sampling, Latin hypercube sampling, and random sampling do not account for the sensitivity of the model’s predictions to input frequencies. Instead, we employ the method of acquisition function [45] to sample input space for the training of complex-valued DeepONet.
@@
4 Computational Experiments
In this section, we demonstrate the performance of the proposed DeepONet architecture in solving Maxwell’s equations. To evaluate its effectiveness across different geometric complexities, we consider two distinct geometries:
  1. 1.
    A three-dimensional sphere with a smooth surface, which serves as a baseline case without singularities.
  2. 2.
    An almond-shaped geometry with sharp corners, which introduces singularities and poses a greater challenge for the neural network to accurately approximate the electromagnetic fields.
This comparison allows us to assess the capability of the proposed DeepONet to handle both smooth and singularity-prone geometries.
4.1 3D Homogeneous Metallic Sphere
Here, we investigate the performance of the proposed complex-valued DeepONet architecture and highlight its advantages over conventional real-valued DeepONets. Specifically, we compare the two approaches in terms of accuracy, convergence, and their ability to capture complex-valued solutions arising from electromagnetic problems at two very large frequencies. This analysis demonstrates the superiority of the complex-valued framework, particularly in handling the inherent coupling between the real and imaginary components of the solution, which is often lost when using real-valued networks.
The computational domain of the problem, shown in Figure 3, consists of a homogeneous metallic sphere with a radius of $r=1\,\text{m}$ (shown as green sphere) and an outer boundary positioned at a distance of $3\lambda$ from the metallic sphere, where $\lambda\,\text{(m)}$ is the wavelength given as
$\lambda=\frac{c}{f}$ (24)
where $c=3\times 10^{8}\,\text{m/s}$ is the speed of light and $f\,\text{(Hz)}$ is the maximum frequency of the electromagnetic wave considered in this problem.
The computational domain is discretized by tetrahedral meshes with $3$ cells per wavelength $(\lambda)$ and is presented in 3(b). Two different problems are considered in this work. For the first study, we consider electromagnetic fields with a maximum frequency of $0.6\,\text{GHz}$. The corresponding wave length $\lambda=0.5~\text{m}$ and the sphere surface resolution is $\frac{1}{6}\,\text{m}$. Under this configuration, the sphere surface is discretized into 1004 triangular facets, the outer boundary is discretized into 1408 triangular facets, and the volume cells are discretized into 25,042 tetrahedral elements with a maximum cell size of $6\lambda$. For a more advanced second study, we consider a much higher maximum frequency of $4.77\,\text{GHz}$, so the corresponding wavelength $\lambda=6.28\times 10^{-2}\,\text{m}$, the sphere surface resolution is $2.10\times 10^{-2}\,\text{m}$, and the sphere is discretized into 137,388 triangular facets. In the second case, we test the scalability of the proposed surrogate model to a significantly larger dataset required to resolve the higher frequency and smaller wavelength.
The training dataset contains the incident fields $\boldsymbol{E}^{i}(x,y,z;f)$ impinging the sphere at different incidence angles, frequencies $(f)$, and spatial locations of sensors, which in this case are coordinates of vertices $(x,y,z)$ of the tetrahedral and triangles. The output of the DeepONet is the total field $\boldsymbol{E}(x,y,z;f)$. To generate the training and testing data, we utilized the Ultra Weak Variational Formulation (UWVF) method [25, 24, 26] to solve the system described by (1). UWVF is a special case of a Trefftz-discontinuous Galerkin (Trefftz-DG) method provided the scattering medium is lossless. The UWVF approach uses an unstructured finite element computational grid and a superposition of plane wave solutions of Maxwell’s equations on each tetrahedral element [40]. UWVF has been found to be efficient for approximating solutions of Maxwell’s equations for a wide range of test problems as shown by Lahivaara et al.[30].
(b) Meshed geometry of the computational domain.
@@ 4 Computational Experiments
To demonstrate the superior performance of the complex-valued DeepONet compared to the vanilla real-valued DeepONet, we used similar architectures for both models, training them on the same 80 frequencies and testing on 21 frequencies sampled from a range of $0.05,0.6]\,\text{GHz}$. The hyperparameters for both models used in our experiments are summarized in Table 1. For the real-valued DeepONet, the real and imaginary parts of a function are trained separately and then combined into a complex number for comparison against the ground-truth data. Note that while the complex-valued DeepONet uses complex parameters (effectively doubling the real parameter count per weight), the real-valued DeepONet requires two separate networks—one for the real part and one for the imaginary part—for each field component. Consequently, the total number of trainable real-valued parameters is approximately the same for both architectures. The training loss histories of the two models are plotted in Figure 4, with the dotted line representing the real-valued DeepONet and the solid line representing the complex-valued DeepONet. The loss stabilizes at around 10,000 epochs, with the complex-valued DeepONet achieving superior training loss across all components of the field. The final testing results are summarized in Figure 4, which shows that the reparametrized complex-valued DeepONet, trained using the two-step method, exhibits the best testing performance. The complex-valued models significantly outperform the vanilla real-valued model by decreasing the error with a factor ranging from 5-7 times. In terms of inference speed, the UWVF numerical solver takes 15 minutes to predict one frequency using 192 cores. In comparison, the complex-valued DeepONet predicts a single frequency in 1.16 seconds on a 24-core CPU or 0.10 seconds on an Nvidia Titan RTX GPU.
@@
Figure 5 shows the correlation between predicted and true total fields. Strong diagonal alignment demonstrates high prediction accuracy for both the electric and magnetic fields. Figure 6 shows the predictions of the total electric and magnetic fields by the complex-valued model at a testing frequency of 0.402 GHz. The 2D visualizations on the sphere surface demonstrate close agreement between predictions and ground truth for both real and imaginary parts of all field components. Figure 7 presents the absolute pointwise errors, which are small throughout the domain. The model is shown to generalize well to the testing frequencies.
@@
Figure 4: Left subfigure shows training loss history for each electromagnetic field using complex-valued DeepONets (solid lines) and real-valued DeepONets (dashed lines), starting from epoch 500 on the logarithmic scale. Each field component is distinguished by color. The figure shows the superior convergence of the complex-valued DeepONet. The right subfigure shows testing performance of DeepONets (the complex-valued DeepONet and the reparametrized complex-valued DeepONet). The plot shows a superior performance of complex-valued DeepONets over real-valued DeepONet.
@@
Figure 6: Comparison between ground truth and predictions of the electromagnetic fields on the sphere geometry at a testing frequency of 0.402 GHz. The top four rows show the real and imaginary parts of the electric field components (top two rows) and magnetic field components (bottom two rows). Each column corresponds to a field component (Ex, Ey, Ez, Hx, Hy, Hz). The predictions agree closely with the ground truth across all components.
Figure 7: Absolute pointwise errors for all six electromagnetic field components on the sphere geometry at a testing frequency of 0.402 GHz. The top row shows errors in the real parts and the bottom row shows errors in the imaginary parts of each component. The errors are small throughout the domain.
In the second example, we tested the capability of the complex-valued DeepONet to handle electromagnetic fields with a much higher maximum frequency of $4.77\,\text{GHz}$. To resolve the higher frequencies and smaller wavelengths, the computational domain is discretized into a finer mesh as described in LABEL:sec:data_generation and the incident field is evaluated on 137,388 sensor locations. To accommodate this increase, a larger DeepONet model is created with a branch network input dimension of 137,388 and the number of neurons per layer raised to 128. The number of hidden layers and other model hyperparameters are unchanged. After training for 100,000 epochs, the training and testing relative $L_{2}$ losses are presented in Table 2. These results show that the complex-valued DeepONet can effectively scale to a much larger dataset to handle high-frequency fields without compromising its accuracy or altering training routine.
Table 2: Relative $L_{2}$ errors for training and testing for complex-valued DeepONet on a much larger dataset with high-frequency fields up to $4.77\,\text{GHz}$
Component Training $L_{2}$ % Testing $L_{2}$ %
@@
4.2 3D Almond shape geometry with sharp corners
To further demonstrate the capability of the complex-valued DeepONet on geometries beyond the simple sphere, we consider the metallic Almond-shaped target. The Almond target is 0.252 m long and features sharp edges and corners along its boundary, making it more challenging than the smooth spherical scatterer. These sharp features are known to be difficult for numerical methods as they can cause field singularities and require higher mesh resolution to capture accurately.
The computational domain is generated based on analytical expressions provided by the ElectroMagnetics Code Consortium (EMCC). Following the mesh parameters used for sphere geomtery in in subsection 4.1, the Almond surface is discretized with a resolution of $\lambda/8$ on the target surface and $\lambda/33$ on the sharp edges. The outer boundary is positioned at $3\lambda$ from the target, where $\lambda$ is the wavelength at the maximum frequency of 1.19 GHz. The incident plane wave impinges on the Almond at an azimuthal angle of $\phi=60^{\circ}$ in the $\theta=90^{\circ}$ plane. The computational domain is discretized into 11,154 tetrahedral elements, resulting in the evaluation of the incident field on 11,154 sensor locations. A three-dimensional view of the meshed Almond geometry is shown in Figure 8.
Figure 8: Three-dimensional view of the Almond geometry with an unstructured tetrahedral mesh. The mesh resolution is $\lambda/8$ on the target surface and $\lambda/33$ on the sharp edges to accurately capture the geometry curvature at 1.19 GHz.
The training and testing dataset are generated using the UWVF method for 61 uniformly sampled incident angles. The dataset is split evenly, with 50% of the data used for training on a uniform grid and 50% for testing. Six separate complex-valued DeepONets are trained for the six components of the electromagnetic field ($E_{x}$, $E_{y}$, $E_{z}$, $H_{x}$, $H_{y}$, $H_{z}$). The network architecture for each component consists of two branch networks with 4 hidden layers of 256 neurons each, and a trunk network with 3 hidden layers of 256 neurons. The branch networks encode the incident field and frequency information into a 200-dimensional latent space, while the trunk network processes spatial coordinates augmented with Fourier features (feature size = 5). All networks use the complex ReLU ($\mathbb{C}\text{ReLU}$) activation function. The models are trained with a single Adam optimizer using a polynomial learning rate schedule, starting at $10^{-3}$ and decaying to $10^{-5}$ over 10,000 steps with a decay power of 0.5. The training is conducted for up to 200,000 epochs for the $E_{x}$ component and 100,000 epochs for all other components.
Note that since the almond geometry is significantly more difficult to predict than the sphere geometry, we employ a larger network and train for more epochs.
The training loss histories for all six field components are shown in Figure 9. The final relative $L_{2}$ errors are summarized in Table 3. The training errors range from 0.15% to 0.71%, while the testing errors range from 0.22% to 1.24%. These low error values indicate that the complex-valued DeepONet generalizes well to unseen incident angles even on this geometrically complex target.
Figure 9: Training and testing loss histories for the complex-valued DeepONet for the 3D Almond shape geometry shown in Figure 8. The loss shows show convergence for all six field components ($E_{x}$, $E_{y}$, $E_{z}$, $H_{x}$, $H_{y}$, $H_{z}$).The training strategy adopted the same early stopping criteria as in the previous experiment to avoid the overfitting.
Table 3: Relative $L_{2}$ errors for training and testing of the complex-valued DeepONet on the Almond geometry at 1.19 GHz
Component Training $L_{2}$ % Testing $L_{2}$ %
$E_{x}$ 0.71 1.24
$E_{y}$ 0.41 0.68
$E_{z}$ 0.16 0.23
$H_{x}$ 0.31 0.37
$H_{y}$ 0.15 0.27
$H_{z}$ 0.31 0.59
Figure 10 presents a comparison between the ground truth and predicted fields for a testing incident angle at $\phi=1^{\circ}$. The top two rows show the real and imaginary parts of the electric field components, while the bottom two rows show the magnetic field components. Figure 11 shows the absolute pointwise errors for all six components. The errors are generally very small across the computational domain, with larger errors concentrated primarily near the sharp edges and corners of the Almond surface.
Figure 10: Comparison between ground truth and predictions of the electromagnetic fields on the Almond geometry shown in Figure 8 for a testing incident angle at $\phi=1^{\circ}$. The top four rows show the real and imaginary parts of the electric field components (top two rows) and magnetic field components (bottom two rows).
Figure 11: Absolute pointwise errors for all six electromagnetic field components on the Almond geometry shown in Figure 8 at a testing incident angle of $\phi=1^{\circ}$. The errors are small throughout the domain, with slightly larger values concentrated near the sharp edges and corners where field singularities occur.
These results demonstrate that the complex-valued DeepONet is capable of predicting electromagnetic fields on geometries with sharp edges and corners. In addition, the model can be scaled up to address more complicated problems where more expressivity is needed
4.3 Acquisition functions for sampling input frequencies
@@
The acquisition function is used to efficiently select training frequencies for the complex-valued DeepONet. We consider the frequency range of $0.05,0.6]\,\text{GHz}$. An initial set of 5 uniformly distributed frequencies in this range are selected for training an ensemble of 3 DeepONets. Then, an additional 15 frequencies are acquired to maximize information gain as determined by the acquisition function. The trained model is subsequently tested on 80 unseen samples.
@@
4.4 DeepONet as surrogate for computing radar cross section (RCS) results
The overarching goal of the proposed DeepONet architecture is to efficiently compute integrated quantities, such as the radar cross section (RCS), which is critical for detecting and characterizing the signatures of different objects in real time. To demonstrate this capability, we compute the RCS using the electromagnetic fields inferred from the proposed DeepONet and compare the results against those obtained from a conventional numerical solver. This comparison allows us to assess both the accuracy and reliability of the DeepONet predictions in capturing essential physical quantities, while highlighting the potential of the approach for real-time applications where rapid evaluation of object signatures is required. Another important motivation for this exercise is to assess the sensitivity of integrated quantities to errors in the inferred fields. Even small inaccuracies in the predicted electromagnetic fields can propagate and result in significantly larger errors in computed integrated quantities such as the radar cross section (RCS). This sensitivity analysis is particularly important because many existing works proposing surrogate models often overlook this aspect, focusing primarily on pointwise field accuracy without evaluating its impact on derived quantities.
@@ 4.4 DeepONet as surrogate for computing radar cross section (RCS) results
The RCS is expressed as the ratio of the power density of the scattered field evaluated in the far-field to the incident power density times the area of a sphere of radius $r$, in the limit $r\to\infty$. The monostatic and bistatic radar cross sections (RCS), computed using field predictions from the complex-valued DeepONet, are presented in Figure 14. For these results, the DeepONet was trained on 60 samples and tested on 40 samples. In Figure 14, the monostatic RCS is calculated at an azimuth angle of $\phi=0$ for frequencies ranging from 0.05 to 0.6 GHz, while the bistatic RCS is shown for a frequency of $0.281$ GHz with $\phi$ varying from 0 to $360$ degrees. For both monostatic and bistatic RCS calculations, the analytical Mie-series solutions are used as the ground truth for comparison. The results of RCS ensures the accuracy proposed DeepONet model for computing the integrated quantity, which is very sensitive to the accuracy of predict electromagnetic field from DeepONets.
Figure 14: Monostatic and bistatic RCS results for the spherical scatterer, shown in Figure 3, computed from DeepONet predictions (orange) and compared to ground-truth Mie-series solutions (blue). The monostatic RCS (left) is evaluated at 40 testing frequencies from the interval $[0.05,0.6]$ GHz. The bistatic RCS (right) is calculated at a single frequency of $0.281$ GHz and plotted against the azimuth angle $\phi$. Predictions (dashed lines with square markers) match the ground truth (solid lines with circle markers) with high accuracy.
The monostatic RCS for the Almond target is presented in Figure 15. The DeepONet was trained on 30 incident angles and tested on 31 unseen angles at a frequency of 1.19 GHz. Figure 15 shows that the predicted RCS values (orange squares) agree well with the ground truth UWVF solutions (blue line).
Figure 15: Monostatic RCS results for the Almond scatterer, shown in Figure 8, at 1.19 GHz as a function of incident azimuth angle $\phi$. The ground truth UWVF solution (blue solid line) is compared with DeepONet predictions (orange squares, 31 testing angles). The predictions agree well with the ground truth across the angular range